# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ 6 Conclusion
In this paper, we have proposed three copula-based metrics for evaluating survival models under dependent censoring. We empirically compared them to existing, independent variants in both synthetic and semi-synthetic datasets. We found that in datasets with adequate instances and censoring, the dependent metrics gave error estimates that were closer to the true error for prediction accuracy, i.e., Brier Score and MAE. We anticipate that others will be able to use these metrics and improve upon our framework.
Acknowledgements.
@@
Appendix A Overview of the appendices
Appendix B provides an overview of the notation. Appendix C describes the main censoring assumptions. Appendix D describes the theory behind copulas. Appendix E describes the existing survival metrics. Appendix F gives details on data generation, preprocessing, survival learners and implementation. Appendix G contains the complete results for the datasets. Appendix H provides derivations of used estimators.
Appendix B Notation
@@
Symbol/Abbr. Definition
@@
$f_{E}$ Probability density function, representing $\text{Pr}(E=t)$
$F_{E}$ Cumulative density function, representing $\text{Pr}(E<t)$
$S_{E}$ Survival function, $S:\mathbb{R}\rightarrow[0,1]$
@@
$h(\cdot\mid\bm{x}_{i})$ Hazard function given the covariates $\bm{x}_{i}$
@@
Table 2: Table of notation.
Appendix C Censoring assumptions
Random censoring. We can define random censoring as follows: “Subjects who are censored at time t should be representative of all the study subjects who remained at risk at time t with respect to their survival experience.” [Kleinbaum and Klein, 2012, Ch. 1]. In other words, we assume that the death rate (or failure rate) of censored subjects is the same as that of uncensored subjects who remain in the risk set. Random censoring is a stronger assumption and more restrictive than what is called independent censoring, which states that censoring is independent within a subgroup.
Definition 5
@@ Appendix C Censoring assumptions
Independent censoring. In the absence of competing risks, we can define independent censoring as follows: “Within any subgroup of interest, subjects who are censored at time $t$ should be representative of all subjects in that subgroup who remained at risk at time $t$ with respect to their survival experience” [Kleinbaum and Klein, 2012, Ch. 1]. In other words, censoring is independent provided that it is random within any subgroup given the covariates.
Definition 6
@@
$h_{A,\text{CE}}(t)$ $=h_{A,\text{NCE}}(t)$ (C.2)
$h_{B,\text{CE}}(t)$ $=h_{B,\text{NCE}}(t)$
@@
The idea behind random and independent censoring is that, even though the censored subjects were not randomly selected, their survival experience would be expected to be the same as if they had been randomly selected from the risk set at time $t$. Consider a clinical trial studying the recovery time of patients after surgery. The patients’ recovery times are measured, but some individuals leave the study early for reasons unrelated to their recovery, such as relocating to another city, personal commitments, or withdrawal for reasons not related to their health (e.g., family obligations). These individuals are censored, but their reason for leaving the study has nothing to do with the outcome (recovery time). Many of the analytic techniques in survival analysis, e.g., the Kaplan-Meier [Kaplan and Meier, 1958] and Nelson-Aalen [Nelson, 1969, Aalen, 1978] population estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [Cox, 1972], rely on the assumption of independent censoring for valid inference in the presence of right-censored data [Kleinbaum and Klein, 2012, Ch. 1]. When censoring is not independent, however, this can introduce bias in survival analysis. Bias arises when censored individuals have a higher likelihood of experiencing the event compared to those who are not censored. Consequently, if a significant portion of censored individuals actually had the event (e.g., died), the estimated survival probability at any time $t$ may overestimate the true survival probability at that time. By example of Kleinbaum and Klein [2012, Ch. 1], consider a drug study where some patients are censored due to the occurrence of side effects. The unobserved survival outcomes of those censored because of side effects may not be representative of those who remain in the study. If patients who experience side effects are more susceptible to the health outcome, e.g., death, assuming independent censoring could lead to an overestimation of their survival rate.
@@
Non-informative censoring. Censoring can be non-informative if the distribution of survival times ($E$) provides no information about the distribution of censoring times ($C$), and vice versa. Otherwise, the censoring is informative. However, we would still need to know which subjects are censored or not censored. The assumption of non-informative censoring is often justifiable when censoring is independent and/or random [Kleinbaum and Klein, 2012, Ch. 1], but cases exist where the censoring is informative but also random and independent: Consider a cohort study, where each time a patient experiences an event (e.g., death), another patient in the study is randomly selected to be censored. This could be a family member who decides to leave the study after an event occurs. If those who are censored are representative of those who remain in the risk set, then the censoring would be random and independent. However, in this case, the censoring would also be informative, as the censoring mechanism would be linked to the time-to-event distribution (since events trigger censoring). In fact, if this were the only mechanism by which individuals were censored, the distribution of survival times would entirely determine the distribution of censoring times, making it highly informative.
Appendix D Copulas and Definitions
@@
We can view bivariate survival analysis as a competing risks problem, where the occurrence of an event (e.g., death) and censoring (e.g., dropout) are mutually-exclusive incidents. Tsiatis [1975] called this the identifiability issue; competing risks data do not allow one to observe two event times simultaneously (one event censors the other), and hence, the data may not identify the dependence structure between event times. Zheng and Klein [1995] gave a partial solution to the identifiability problem of dependent censoring by an assumed copula between two event times. A copula is a function that links two or more random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. A mathematician, Abe Sklar, first used the word “copula” in his study of probabilistic metric space [Sklar, 1959]. In his paper, he gave a mathematical definition of copulas and established the most fundamental theorem about copulas, known as Sklar’s theorem [Sklar, 1959]. It demonstrated that any joint cumulative density can be written in terms of a copula over the quantiles of its marginal cumulative densities. Among the various types of copulas, the Archimedean copula is the most used, and enables the modeling of dependencies in arbitrarily high dimensions using a single parameter. This family of copulas is defined by a convex generator function, $\varphi:[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3]. Given event time $E$ and censoring time $C$, we can define an Archimedean copula as:
$\begin{split}P(E>t,C>c)=\varphi^{-1}\big{[}\varphi\{S_{E}(t)\}+\varphi\{S_{C}(t)\}\big{]},\end{split}$ (D.1)
@@
where $\varphi$ is the generator function, and $S_{E}(t)=P(E>t)$ and $S_{C}(t)=P(C>t)$ are the marginal survival functions. In the bivariate case, $\varphi$ describes the dependency structure (i.e., correlation) between two random variables, e.g.$E$ and $C$, and the (rank) correlation can described with Kendall’s tau ($\tau$):
$\begin{split}\tau=&\quad\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})\geq 0\}-\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})<0\},\end{split}$ (D.2)
@@
where ($e_{1}$, $c_{1}$) and ($e_{2}$, $c_{2}$) are sampled from a (survival) copula. The Clayton [Clayton, 1978] and Frank [Frank, 1979] copulas are two common Archimedean copulas in survival literature. Within these families, the copula $C_{\theta}$ is parameterized by $\theta$, interpreted as the degree of dependence between ${E}$ and ${C}$ in the bivariate case. When $\theta\approx 0$, it yields the independence copula, thus a larger value of $\theta$ implies greater dependency between the marginal distributions. Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities and the following properties Nelsen [2006]:
@@ 2. Uniform Margins: For all $i\in\{1,\dots,m\}$, if $\forall j\neq i\Rightarrow u_{j}=1$, then
Let ${E}$ be the event distribution, ${C}$ the censoring distribution, $\bm{X}$ a vector of covariates, and $S_{E\mid\bm{X}}(t\mid\bm{X})=\text{Pr}(E>t\mid\bm{X})$ and $S_{C\mid\bm{X}}(t\mid\bm{X})=\text{Pr}(C>t\mid\bm{X})$ the marginal survival functions given $\bm{X}$. We can then define a bivariate (survival) copula $C_{\theta}$ [Nelsen, 2006] that describes the degree of dependence between $E$ and $C$ as follows:
$\text{Pr}(E>t,C>t\mid\bm{X})\ =\ C_{\theta}\{\,S_{E\mid\bm{X}}(t\mid\bm{X}),\,S_{C\mid\bm{X}}(t\mid\bm{X})\,\}.$ (D.3)
Kendall’s tau ($\tau$) can be similarly defined as:
$\begin{split}\tau=&\quad\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})\geq 0\mid\bm{X}\ \}-\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})<0\mid\bm{X}\ \},\end{split}$ (D.4)
where ($e_{1}$, $c_{1}$) and ($e_{2}$, $c_{2}$) are sampled from the model (D.3). Kendall’s $\tau$ can be solely expressed as a function of the copula $C_{\theta}$ by [Emura and Chen, 2018, Ch. 3]:
$\tau_{\theta}=4\int_{0}^{1}\int_{0}^{1}C_{\theta}(u,v)\,C_{\theta}(du,dv),$ (D.5)
which implies that Kendall’s $\tau$ does not depend on the marginal survival functions. An Archimedean copula is a type of copula that is defined using a generator function, which allows for a flexible way to model dependence structures between random variables, e.g.$E$ and $C$. Such a copula is defined as [Emura and Chen, 2018, Ch. 3]:
$C_{\theta}(u,v)\ =\ \varphi_{\theta}^{-1}\left(\,\varphi_{\theta}(u)+\varphi_{\theta}(v)\,\right),$ (D.6)
@@
Appendix E Existing Evaluation Metrics
E.1 Concordance Index
The Harrell’s concordance index (CI) quantifies the proportion of concordant pairs among all comparable pairs [Harrell Jr et al., 1996]. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, CI is computed as:
$\text{CI}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.$
Some definitions of CI [Hartman et al., 2023] incorporate tie-breaking rules for cases where risk scores are identical. Specifically, a score of 0.5 may be assigned to tied pairs in the numerator. However, for simplicity, we omit ties in our discussion.
As shown in Uno et al. [2011], this calculation converges to a censoring-dependent probability:
@@ E.1 Concordance Index
$\text{CI}_{\text{IPCW}}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},$ (E.1)
where $\hat{G}(t)$ represents the marginal survival probability for censoring at time $t$, typically estimated using the Kaplan-Meier (KM) estimator. The inverse probability censoring weight (IPCW) is given by $1/\hat{G}(t)$. In cases where censoring probability depends on features (i.e., $\hat{G}(t\mid\bm{x}_{i})$), a survival model such as CoxPH can be used for estimation. These two metrics are implemented in package SurvivalEVAL [Qi et al., 2023b].
E.2 Integrated Brier Score
The Brier Score (BS) quantifies the accuracy of probabilistic survival predictions at a specific time $t^{*}$ [Graf et al., 1999]. It measures the squared error between the predicted survival probability and the observed survival status at $t^{*}$:
To incorporate the censored subjects in the calculation, BS also requires estimating the IPCW to uniformly transfer a censored subject’s weight (originally 1) to subjects with known status at that time [Vock et al., 2016]. Formally:
$\text{BS}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}(\cdot),\hat{G}(\cdot),t^{*}\right)\ =\ \frac{1}{N}\sum_{i}\left(\frac{\delta_{i}\cdot\mathbbm{1}[t_{i}\leq t^{*}]\cdot\hat{S}(t^{*}\mid\bm{x}_{i})^{2}}{\hat{G}(t_{i})}+\frac{\mathbbm{1}[t_{i}>t^{*}]\cdot(1-\hat{S}(t^{*}\mid\bm{x}_{i}))^{2}}{G(t^{*})}\right),$

### v3

@@ 7 Conclusion
In this paper, we have proposed three copula-based metrics for evaluating survival models under dependent censoring. We empirically compared them to existing, independent variants in both synthetic and semi-synthetic datasets. We found through empirical evaluation that our metrics can more accurately estimate model error than existing alternatives, though their effectiveness depends on the dependence structure in the data and the amount of censoring. We anticipate that others will be able to use these metrics and improve upon our framework.
Acknowledgments and Disclosure of Funding
@@
Appendix A Notation
@@
Table 3: Table of notation.
Symbol Definition
@@
$f_{T}$ Probability density function, representing $\text{Pr}(T=t)$
$F_{T}$ Cumulative density function, representing $\text{Pr}(T<t)$
$S_{T}$ Survival function, $S:\mathbb{R}\rightarrow[0,1]$
@@
$h(\cdot\mid\bm{x})$ Hazard function given the covariates $\bm{x}$
@@
Appendix B Censoring assumptions
Random censoring. We define random censoring as follows: “Subjects who are censored at time t should be representative of all the study subjects who remained at risk at time t with respect to their survival experience.” [5, Ch. 1]. In other words, we assume that the death rate (or failure rate) of censored subjects is the same as that of uncensored subjects who remain in the risk set. Random censoring is a stronger assumption and more restrictive than what is called independent censoring, which states that censoring is independent within a subgroup.
Definition 3.
@@ Appendix B Censoring assumptions
Independent censoring. In the absence of competing risks, we define independent censoring as follows: “Within any subgroup of interest, subjects who are censored at time $t$ should be representative of all subjects in that subgroup who remained at risk at time $t$ with respect to their survival experience” [5, Ch. 1]. In other words, censoring is independent provided that it is random within any subgroup given the covariates.
Definition 4.
@@
$h_{A,\text{CE}}(t)$ $=h_{A,\text{NCE}}(t),$ (B.2)
$h_{B,\text{CE}}(t)$ $=h_{B,\text{NCE}}(t),$
@@
The idea behind random and independent censoring is that, even though the censored subjects were not randomly selected, their survival experience would be expected to be the same as if they had been randomly selected from the risk set at time $t$. Consider a clinical trial studying the recovery time of patients after surgery. The patients’ recovery times are measured, but some individuals leave the study early for reasons unrelated to their recovery, such as relocating to another city, personal commitments, or withdrawal for reasons not related to their health (e.g., family obligations). These individuals are censored, but their reason for leaving the study has nothing to do with the outcome (recovery time). Many of the analytic techniques in survival analysis, e.g., the Kaplan-Meier [6] and Nelson-Aalen [7, 8] population estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [9], rely on the assumption of independent censoring for valid inference in the presence of right-censored data [5, Ch. 1]. When censoring is not independent, however, this can introduce bias in the analysis. Such bias can arise when censored individuals have a higher likelihood of experiencing the event compared to those who are not censored. Consequently, if a significant portion of censored individuals had the event (e.g., died), the estimated survival probability at any time point may overestimate the true survival probability at that time. By example of [5, Ch. 1], consider a drug study where some patients are censored due to the occurrence of side effects. The unobserved survival outcomes of those censored because of side effects may not be representative of those who remain in the study. If patients who experience side effects are more susceptible to the health outcome, e.g., death, assuming independent censoring could lead to an overestimation of their survival rate.
@@
Non-informative censoring. Censoring can be non-informative if the distribution of survival times ($T$) provides no information about the distribution of censoring times ($C$), and vice versa. Otherwise, the censoring is informative. However, we would still need to know which subjects are censored or not. The assumption of non-informative censoring is often justifiable when censoring is independent and/or random [5, Ch. 1], but cases exist where the censoring is informative but also random and independent: Consider a cohort study, where each time a patient experiences an event (e.g., death), another patient in the study is randomly selected to be censored. This could be a family member who decides to leave the study after an event occurs. If those who are censored are representative of those who remain in the risk set, then the censoring would be random and independent. However, in this case, the censoring would also be informative, as the censoring mechanism would be linked to the time-to-event distribution (since events trigger censoring). In fact, if this were the only mechanism by which individuals were censored, the distribution of survival times would entirely determine the distribution of censoring times, making it highly informative.
Appendix C Copulas and definitions
@@
We can view bivariate survival analysis as a competing risks problem, where the occurrence of an event (e.g., death) and censoring (e.g., dropout) are mutually-exclusive incidents. Tsiatis [40] referred to this as the identifiability issue; competing risks data do not allow one to observe two event times simultaneously (one event censors the other), and hence, the data may not identify the dependence structure between event times. Zheng and Klein [12] gave a partial solution to the identifiability problem of dependent censoring by an assumed copula between two event times. A copula is a function that links two or more random variables by specifying their dependence structure [4, Ch. 1]. A mathematician, Abe Sklar, first used the word “copula” in his study of probabilistic metric space [42]. In his paper, he gave a mathematical definition of copulas and established the most fundamental theorem about copulas, known as Sklar’s theorem [42]. It demonstrated that any joint cumulative density can be written in terms of a copula over the quantiles of its marginal cumulative densities. Among the various types of copulas, the Archimedean copula is the most used, and enables the modeling of dependencies in arbitrarily high dimensions using a single parameter. This family of copulas is defined by a convex generator function, $\varphi:[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [4, Ch. 3]. Given event time $T$ and censoring time $C$, we can define an Archimedean copula as:
$\begin{split}P(T>t,C>c)=\varphi^{-1}\big{[}\varphi\{S_{T}(t)\}+\varphi\{S_{C}(c)\}\big{]},\end{split}$ (C.1)
@@
where $\varphi$ is the generator function, and $S_{T}(t)=P(T>t)$ and $S_{C}(t)=P(C>c)$ are the marginal survival functions. In the bivariate case, $\varphi$ describes the dependence structure (i.e., correlation) between two random variables, e.g.$T$ and $C$, and the (rank) correlation can be described with Kendall’s tau ($\tau$):
$\begin{split}\tau=&\quad\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})\geq 0\}-\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})<0\},\end{split}$ (C.2)
@@
where ($t_{1}$, $c_{1}$) and ($t_{2}$, $c_{2}$) are sampled from a (survival) copula. The Clayton [10] and Frank [24] copulas are two common Archimedean copulas in survival literature. Within these families, the copula $C_{\theta}$ is parameterized by $\theta$, interpreted as the degree of dependence between ${T}$ and ${C}$ in the bivariate case. When $\theta\approx 0$, it yields the independence copula; thus, a larger value of $\theta$ implies greater dependency between the marginal distributions. Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities and the following properties [43]:
@@ 2. Uniform Margins: For all $i\in\{1,\dots,m\}$, if $\forall j\neq i\Rightarrow u_{j}=1$, then
Let $T$ be the event distribution, $C$ the censoring distribution, $\bm{x}$ a vector of covariates, and $S_{T\mid\bm{x}}(t\mid\bm{x})=\text{Pr}(T>t\mid\bm{x})$ and $S_{C\mid\bm{x}}(t\mid\bm{x})=\text{Pr}(C>c\mid\bm{x})$ the marginal survival functions given $\bm{x}$. We can then define a bivariate (survival) copula $C_{\theta}$ [43] that describes the degree of dependence between $T$ and $C$ as follows:
$\text{Pr}(T>t,C>c\mid\bm{x})\ =\ C_{\theta}\{\,S_{T}(t\mid\bm{x}),\,S_{C}(c\mid\bm{x})\,\}.$ (C.3)
Kendall’s tau ($\tau$) can be defined as:
$\begin{split}\tau=&\quad\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})\geq 0\mid\bm{x}\ \}-\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})<0\mid\bm{x}\ \},\end{split}$ (C.4)
where ($t_{1}$, $c_{1}$) and ($t_{2}$, $c_{2}$) are sampled from the model (C.3). Kendall’s $\tau$ can be solely expressed as a function of the copula $C_{\theta}$ by [4, Ch. 3]:
$\tau_{\theta}=4\int_{0}^{1}\int_{0}^{1}C_{\theta}(u,v)\,C_{\theta}(du,dv),$ (C.5)
which implies that Kendall’s $\tau$ does not depend on the marginal survival functions. An Archimedean copula is a type of copula that is defined using a generator function, which allows for a flexible way to model dependence structures between random variables, e.g.$T$ and $C$. Such a copula is defined as [4, Ch. 3]:
$C_{\theta}(u,v)\ =\ \varphi_{\theta}^{-1}\left(\,\varphi_{\theta}(u)+\varphi_{\theta}(v)\,\right),$ (C.6)
@@
Appendix D Existing evaluation metrics
D.1 Concordance Index
Harrell’s concordance index (CI) quantifies the proportion of concordant pairs among all comparable pairs [29]. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, CI is computed as:
$\text{CI}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.$ (D.1)
Some definitions of CI [44] incorporate tie-breaking rules for cases where risk scores are identical. Specifically, a score of 0.5 may be assigned to tied pairs in the numerator. However, for simplicity, we omit ties in our discussion.
As shown in [30], this calculation converges to a censoring-dependent probability:
@@ D.1 Concordance Index
$\text{CI}_{\text{IPCW}}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},$ (D.2)
where $\hat{G}(t)$ represents the marginal survival probability for censoring at time $t$, typically estimated using the Kaplan-Meier (KM) estimator. The inverse probability censoring weight (IPCW) is given by $1/\hat{G}(t)$. In cases where censoring probability depends on features (i.e., $\hat{G}(t\mid\bm{x}_{i})$), a survival model such as CoxPH can be used for estimation. These two metrics are implemented in the SurvivalEVAL package [45].
D.2 Integrated Brier Score
The Brier Score (BS) quantifies the accuracy of probabilistic survival predictions at a specific time $t^{*}$ [32]. It measures the squared error between the predicted survival probability and the observed survival status at $t^{*}$:
To incorporate the censored subjects in the calculation, BS also requires estimating the IPCW to uniformly transfer a censored subject’s weight (originally 1) to subjects with known status at that time [46]. Formally:
BS ($\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}_{T}(\cdot\mid\cdot),\hat{G}(\cdot),t^{*}\right)\ =$
$\ \frac{1}{N}\sum_{i}\left(\frac{\delta_{i}\cdot\mathbbm{1}[t_{i}\leq t^{*}]\cdot\hat{S}_{T}(t^{*}\mid\bm{x}_{i})^{2}}{\hat{G}(t_{i})}+\frac{\mathbbm{1}[t_{i}>t^{*}]\cdot(1-\hat{S}_{T}(t^{*}\mid\bm{x}_{i}))^{2}}{G(t^{*})}\right).$ (D.3)